# The role of the exchange-Coulomb potential in two-dimensional electron transport

### v1

@@ I Introduction
Most theoretical treatments of exchange in low-dimensional electron systems adopt either a static or a weakly dynamical approach. In equilibrium, exchange is commonly incorporated through density-functional or Thomas–Fermi–Dirac type approximations that modify the equation of state and the static screening length Kohn and Sham (1965); Dirac (1930); Perdew and Zunger (1981). In dynamical problems, such as plasmons propagation and Coulomb drag instabilities, exchange and correlations are usually included via local-field factors that renormalise the RPA susceptibility and the effective interaction potential Calmels and Gold (1995); Richardson and Ashcroft (1994), or through spin-dependent extensions for spin Coulomb drag D’Amico and Vignale (2003); Weber et al. (2005). These approaches are tailored to near-equilibrium and weakly perturbed regimes, and incorporate exchange only through static renormalisations of the plasmon dispersion or equilibrium response functions. As a result, the exchange field does not enter the kinetic equation as a genuinely nonlocal, momentum-dependent force and does not evolve self-consistently with the full nonequilibrium distribution function Brodin et al. (2019); Zamanian et al. (2015). Consequently, such models are not suited to address situations where exchange modifies the local acceleration, phase-space flow, and nonlinear evolution of an electron gas subject to strong spatial inhomogeneities or driven far from equilibrium Haas (2022); Manfredi et al. (2021); Yadav and Kumar (2025); Kitayev and Dubinov (2024).
@@ II Quantum model for the 2D electron gas
with $V_{\mathbf{q}_{\perp}}^{\text{(2D)}}=e^{2}/(2\varepsilon\mathcal{A}q_{\perp})$ being the two-dimensional Fourier transform of $V(\mathbf{r}_{\perp})$. The lowest-subband energy can be set to zero without loss of generality.
Next, we apply the Hartree-Fock decomposition to the Coulomb interaction in Eq. (5), as described in Ref. Figueiredo et al. (2024). In practice, this procedure amounts to neglect higher-order quantum fluctuations of fermionic operators by keeping only the mean-field contributions plus a small correction associated with electron–elecron collisions. We are led to the effective Hamiltonian
@@
In what follows we assume that electron-electron correlations are small and drop $\hat{\mathcal{C}}$. As a result, Eqs. (6) and (7) establish a closed equation for the electron Wigner function $f_{\mathbf{k}}(\mathbf{r})\equiv\sum_{\mathbf{k}^{\prime}}e^{i\mathbf{k}^{\prime}\cdot\mathbf{r}}\langle\hat{c}_{\mathbf{k}-\mathbf{k}^{\prime}/2}^{\dagger}\hat{c}_{\mathbf{k}+\mathbf{k}^{\prime}/2}\rangle$ Figueiredo et al. (2024), which reads
@@
Above, $\mathcal{H}_{\mathbf{k}}(\mathbf{r})=\sum_{\mathbf{k}^{\prime}}e^{i\mathbf{r}\cdot\mathbf{k}^{\prime}}\mathcal{H}_{\mathbf{k}-\mathbf{k}^{\prime}/2,\mathbf{k}+\mathbf{k}^{\prime}/2}$ is the Wigner transform of $\mathcal{H}_{\mathbf{k},\mathbf{k}^{\prime}}$, given by
@@ III Fluid model with exchange corrections
Moreover $P$ is the usual kinetic pressure.
@@ III.2 Degenerate plasmon dispersion
An immediate consequence of the negative exchange contribution $\widetilde{v}$ is that nonzero imaginary parts of the frequency are now possible even for a single-sheet configuration if $U^{2}<0$, which requires $n_{0}\lesssim 10^{14}\,\text{cm}^{-2}$ for typical values of the effective mass. Denoting $\omega(\mathbf{k})=\Omega(\mathbf{k})+i\gamma(\mathbf{k})$ and assuming that $U^{2}$ is negative, we get a finite value of $\gamma$ for $k>k_{c}$, with $k_{c}=\alpha/|U^{2}|$ the lowest unstable mode. For small displacements $\delta k=k-k_{c}$, it follows that
@@
Figure 2: Real (a) and imaginary (b) parts of the plasmon dispersion including the effect of exchange (solid curves) and neglecting exchange (dashed curve) for varying densities $n_{0}=10^{10}\,\text{cm}^{-2}$ (orange) and $n_{0}=10^{11}\,\text{cm}^{-2}$ (blue) with equilibrium velocity $u_{0}=5\times 10^{4}\,\text{m}\text{s}^{-1}$. (c): real part of the plasmon dispersion for growing values of the density. The solid curve corresponds to the critical density ($n_{c}\sim 2\times 10^{14}\,\text{cm}^{-2}$) above which the dispersion is real. ($m=0.067m_{e}$)
@@ IV.1 Isotropic equilibrium
In the case of zero macroscopic current and equal Fermi levels, the roots admit the following solution
@@
For sufficiently small $d$, these reduce to optical $\omega_{+}\simeq\pm\sqrt{\alpha k}$ and acoustic $\omega_{-}\simeq\pm\sqrt{U^{2}}k$ modes. Clearly if $U^{2}>0$ no instability can occur for $\mathbf{u}_{0,i}=0$. Remarkably, in the presence of sufficiently strong exchange, the acoustic branch becomes unstable even at $\mathbf{u}_{0,i}=0$, with growth rate $\gamma=|U|k$, which is different from $\gamma=\sqrt{\alpha k}$ obtained in Eq. (26) for isolated plates. For typical values, we can verify that $|U|k\gg\sqrt{\alpha k}$, which means that the electrostatic interaction between the two fluids leads to an increase of the growth rate even in the absence of streaming as long as $d>1/k_{F}\sim\text{nm}$ and $n<10^{14}\,\text{cm}^{-2}$.
@@ IV.2 Finite equilibrium currents
Figure 3: Acoustic and optical plasmons (a) if exchange is neglected and (b) if exchange is considered, the latter displaying acoustic-optical mode coupling ($n_{0,2}/n_{0,1}=0.1$, $u_{0,2}/u_{0,1}=0.5$, $d=1\,\textup{\AA }$).
@@ V.2 Exchange-mediated dynamical screening
The developed kinetic framework can also be used to study how exchange modifies the dynamical screening of a localized impurity in a degenerate 2DEG. This problem is directly relevant for semiconductor heterostructures Nixon and Davies (1990), where remote dopants, interface roughness and narrow gates generate inhomogeneous electrostatic landscapes that are screened by the mobile carriers Borisov et al. (2004).
@@
Figure 9: Space-time evolution of the density perturbation $\delta n(x,t)/n_{0}$ in the Hartree (left panel) and Hartree–Fock (right panel) cases. The simulation parameters are those of Fig. 8. For the classical evolution, density perturbations disperse away from the impurity, while in the Hartree–Fock case the density remains localized and deepens over time due to exchange effects. At the final times, Friedel-like oscillations are visible around the impurity in the Hartree–Fock case, due to the excitation of small wavelenghts close of the order of $k_{F}^{-1}$.
@@ V.4.1 Drag force from the kinetic equation
To formulate the problem, we generalise the Vlasov equation to a bilayer geometry. The Wigner function $f^{(\ell)}_{k}(x,t)$ for layer $\ell=1,2$ obeys
@@
where the left-hand side is that of Eq. (30) to which we add an external drive $F_{\ell}^{\mathrm{ext}}=-eE_{\ell}^{\text{ext}}$. Moreover, for sufficiently small $d$ (typically of order $\sim 1/k_{F}$ or smaller), interlayer Coulomb collisions $\mathcal{C}_{\ell,\ell^{\prime}}=\delta_{\ell^{\prime},\overline{\ell}}\,\mathcal{I}_{\ell,\overline{\ell}}$ are not negligible and are known to contribute to the drag signal substancially Chen et al. (2015). The form of the interlayer collision integral for Coulomb-drag configurations has been derived in Ref. Messica and Gutman (2024) in the Born approximation, and reduces to a frictional force of the form

### v2

@@ I Introduction
Most theoretical treatments of exchange in low-dimensional electron systems rely on either static or weakly dynamical approximations. For equilibrium problems, exchange is typically incorporated through density-functional or Thomas–Fermi–Dirac type schemes, which renormalise the equation of state and the static screening length Kohn and Sham (1965); Dirac (1930); Perdew and Zunger (1981). In dynamical settings, including plasmon propagation and Coulomb-drag configurations, exchange and correlations are most often accounted for through local-field factors that renormalise the RPA susceptibility and the effective interaction Calmels and Gold (1995); Richardson and Ashcroft (1994), or via spin-resolved extensions developed in the context of spin Coulomb drag D’Amico and Vignale (2003); Weber et al. (2005). These frameworks are inherently restricted to near-equilibrium and weakly perturbed regimes, as exchange enters only through static renormalisations of response functions or collective-mode dispersions. As a result, the exchange field does not appear as a genuinely nonlocal, momentum-dependent force in the kinetic equation, nor does it evolve self-consistently with the full nonequilibrium distribution function Brodin et al. (2019); Zamanian et al. (2015). This limitation prevents such approaches from capturing regimes in which exchange directly alters local acceleration, phase-space transport, and nonlinear dynamics in strongly inhomogeneous or far-from-equilibrium electron gases Haas (2022); Manfredi et al. (2021); Yadav and Kumar (2025); Kitayev and Dubinov (2024).
@@ II Quantum model for the 2D electron gas
with $V_{\mathbf{q}_{\perp}}^{\text{(2D)}}=e^{2}/(2\varepsilon\mathcal{A}q_{\perp})$ being the two-dimensional Fourier transform of $V(\mathbf{r}_{\perp})$. The lowest-subband energy can be set to zero without loss of generality. Moreover, we consider a nonmagnetized, unpolarized 2DEG governed by the spin-independent Coulomb interaction of Eq. (5) and neglect any relativistic correction such as Zeeman splitting or spin-orbit coupling. In this situation the dynamics is identical for each spin component, so for notational simplicity we have suppressed the spin index in the intermediate steps.
Next, we apply the Hartree-Fock decomposition to the Coulomb interaction in Eq. (5), as described in Ref. Figueiredo et al. (2024). In practice, this procedure amounts to neglect higher-order quantum fluctuations of fermionic operators by keeping only the mean-field contributions plus a small correction associated with electron–electron collisions. We are led to the effective Hamiltonian
@@
In what follows we restrict attention to regimes where electron-electron correlations beyond mean field are subleading and neglect the collision operator $\hat{\mathcal{C}}$. This approximation is justified for dilute, low-temperature 2DEGs in which the characteristic plasmon and drift frequencies exceed the electron-electron scattering rate, so that the dynamics is dominated by long-wavelength, nearly collisionless collective motion Bonitz (2016). Under these conditions, Eqs. (6) and (7) define a closed evolution equation for the electronic Wigner function $f_{\mathbf{k}}(\mathbf{r})\equiv\sum_{\mathbf{k}^{\prime}}e^{i\mathbf{k}^{\prime}\cdot\mathbf{r}}\langle\hat{c}_{\mathbf{k}-\mathbf{k}^{\prime}/2}^{\dagger}\hat{c}_{\mathbf{k}+\mathbf{k}^{\prime}/2}\rangle$ Figueiredo et al. (2024), which reads
@@
In particular, we use the spin-summed Wigner distribution $f_{\mathbf{k}}(\mathbf{r},t)=\sum_{\sigma=\uparrow,\downarrow}f_{\mathbf{k}\sigma}(\mathbf{r},t)$ so that phase-space moments of $f_{\mathbf{k}}$ directly yield the physical density and current of an unpolarized 2DEG. Extensions to the polarized/magnetized case would require evolving the spin-resolved distributions $f_{\mathbf{k}\uparrow}$ and $f_{\mathbf{k}\downarrow}$ and adding the corresponding spin-dependent terms to the Hamiltonian.
In Eq. (10), $\mathcal{H}_{\mathbf{k}}(\mathbf{r})=\sum_{\mathbf{k}^{\prime}}e^{i\mathbf{r}\cdot\mathbf{k}^{\prime}}\mathcal{H}_{\mathbf{k}-\mathbf{k}^{\prime}/2,\mathbf{k}+\mathbf{k}^{\prime}/2}$ is the Wigner transform of $\mathcal{H}_{\mathbf{k},\mathbf{k}^{\prime}}$, given by
@@ III Fluid model with exchange corrections
Moreover $P$ is the usual kinetic pressure,
$P(\mathbf{r},t)=\frac{1}{\mathcal{A}}\sum_{\mathbf{k}}\frac{\hbar\mathbf{k}\otimes\hbar\mathbf{k}}{m}f_{\mathbf{k}}(\mathbf{r}).$ (18)
@@ III.2 Degenerate plasmon dispersion
An immediate consequence of the negative exchange contribution $\widetilde{v}$ is that nonzero imaginary parts of the frequency are now possible even for a single-sheet configuration if $U^{2}<0$, which requires $n_{0}\lesssim 10^{14}\,\text{cm}^{-2}$ for typical GaAs effective mass $m=0.067m_{e}$. Denoting $\omega(\mathbf{k})=\Omega(\mathbf{k})+i\gamma(\mathbf{k})$ and assuming that $U^{2}$ is negative, we get a finite value of $\gamma$ for $k>k_{c}$, with $k_{c}=\alpha/|U^{2}|$ the lowest unstable mode. For small displacements $\delta k=k-k_{c}$, it follows that
@@
Figure 2: Real (a) and imaginary (b) parts of the plasmon dispersion including the effect of exchange (solid curves) and neglecting exchange (dashed curve) for varying densities $n_{0}=10^{10}\,\text{cm}^{-2}$ (orange) and $n_{0}=10^{11}\,\text{cm}^{-2}$ (blue) with equilibrium velocity $u_{0}=5\times 10^{4}\,\text{m}\text{s}^{-1}$. (c): real part of the plasmon dispersion for growing values of the density. The solid curve corresponds to the critical density ($n_{c}\sim 2\times 10^{14}\,\text{cm}^{-2}$) above which the dispersion is real. (typical GaAs parameter $m=0.067m_{e}$)
@@ IV.1 Isotropic equilibrium
In the case of zero macroscopic current and equal Fermi levels, the roots of Eq. (29) admit the following solution
@@
For sufficiently small $d$, these reduce to optical $\omega_{+}\simeq\pm\sqrt{\alpha k}$ and acoustic $\omega_{-}\simeq\pm\sqrt{U^{2}}k$ modes. Clearly if $U^{2}>0$ no instability can occur for $\mathbf{u}_{0,i}=0$. Remarkably, in the presence of sufficiently strong exchange, the acoustic branch becomes unstable even at $\mathbf{u}_{0,i}=0$, with growth rate $\gamma=|U|k$, which is different from $\gamma=\sqrt{\alpha k}$ obtained in Eq. (27) for isolated plates. For typical values, we can verify that $|U|k\gg\sqrt{\alpha k}$, which means that the electrostatic interaction between the two fluids leads to an increase of the growth rate even in the absence of streaming as long as $d>1/k_{F}\sim\text{nm}$ and $n<10^{14}\,\text{cm}^{-2}$ for typical GaAs parameters.
@@ IV.2 Finite equilibrium currents
Figure 3: Acoustic and optical plasmons (a) if exchange is neglected and (b) if exchange is considered, the latter displaying acoustic-optical mode coupling ($n_{0,2}/n_{0,1}=0.1$, $u_{0,2}/u_{0,1}=0.5$, d=10d=10\,nm.)
@@ V.2 Exchange-mediated dynamical screening
The developed kinetic framework can also be used to study how exchange modifies the dynamical screening of a localized impurity in a degenerate 2DEG. This problem is directly relevant for semiconductor heterostructures Nixon and Davies (1990), where remote dopants, interface roughness and narrow gates generate inhomogeneous electrostatic landscapes that are screened by the mobile carriers Borisov et al. (2004). Most treatments of exchange in impurity screening are formulated in equilibrium or within linear response, where exchange is incorporated either via static energy functionals (renormalising the compressibility/Thomas–Fermi scale) or via local-field factors that dress the RPA susceptibility and screened interaction Grinberg and Luryi (1987); Asgari et al. (2006); Singwi et al. (1968). These approaches are naturally expressed in $(q,\omega)$ space around an equilibrium state, so exchange enters as a renormalisation of response functions or mode dispersions. Here, instead, screening is obtained directly in real space and real time by evolving the Hartree–Fock–Wigner equation after a localised perturbation, giving the transient build-up of $\delta n(x,t)$ and the corresponding force profile without assuming an a priori dielectric function.
@@
Figure 9: Space-time evolution of the density perturbation $\delta n(x,t)/n_{0}$ in the Hartree (left panel) and Hartree–Fock (right panel) cases. The simulation parameters are those of Fig. 8. For the classical evolution, density perturbations disperse away from the impurity, while in the Hartree–Fock case the density remains localized and deepens over time due to exchange effects. At the final times, Friedel-like oscillations are visible around the impurity in the Hartree–Fock case, due to the excitation of small wavelengths close of the order of $k_{F}^{-1}$.
@@ V.4.1 Drag force from the kinetic equation
To formulate Coulomb drag in coupled layers, we generalise the exchange-corrected Vlasov equation introduced above by including both an external drive and interlayer collision processes. The Wigner function $f^{(\ell)}_{k}(x,t)$ for layer $\ell=1,2$ obeys
@@
where the left-hand side is that of Eq. (31) to which we add an external drive $F_{\ell}^{\mathrm{ext}}=-eE_{\ell}^{\text{ext}}$. Moreover, for sufficiently small $d$ (typically of order $\sim 1/k_{F}$ or smaller), interlayer Coulomb collisions $\mathcal{C}_{\ell,\ell^{\prime}}=\delta_{\ell^{\prime},\overline{\ell}}\,\mathcal{I}_{\ell,\overline{\ell}}$ are not negligible and are known to contribute to the drag signal substantially Chen et al. (2015). The form of the interlayer collision integral for Coulomb-drag configurations has been derived in Ref. Messica and Gutman (2024) in the Born approximation, and reduces to a frictional force of the form